A gas turbine receives 1 lb/sec air at 300 atm, 450°R and expands it reversibly to 30 atm in an adiabatic steady flow manner; AP = 0; AK = 0. (a) Use the deviation charts and find the exit air temperature and the work. (b) Solve using ideal gas theory and compare with foregoing answers.


1
Expert's answer
2021-12-15T16:22:26-0500

"\frac{p_1}{T_1}=\frac{p_2}{T_2},"

"T_2=T_1\frac{p_2}{p_1}=45°R,"

"\dot A=\dot \nu R \Delta T=1.9~\frac{kJ}s."